OpenStudy (boxman61):

Solving Rational Equations- 4/4x-1 - 1/x-1 = 2/4x-1 so far i have.. (4x-1)(4/4x-1) - (4x-1)(1/x-1) = (4x-1)(2/4x-1) 4 - (4)(1) = 2 4-4=2 0=2 Answer is -1/2 just no idea how to get there

OpenStudy (boxman61):

correct answer is -1/2 Teacher gives question and answer

OpenStudy (anonymous):

I'll work it out and get back to you.

OpenStudy (boxman61):

thanks

OpenStudy (phi):

so far i have.. (4x-1)(4/4x-1) - (4x-1)(1/x-1) = (4x-1)(2/4x-1) That is not a bad start. notice that you can cancel things \[\cancel{(4x-1)}\frac{4}{\cancel{4x-1}}- \frac{4x-1}{x-1}= \cancel{(4x-1)}\frac{2}{\cancel{4x-1}}\]

OpenStudy (phi):

because (4x-1)/(4x-1) = 1

OpenStudy (phi):

can you solve \[4- \frac{4x-1}{x-1}=2\]

OpenStudy (boxman61):

ahh feel like an idiot lol...thanks everybody

OpenStudy (anonymous):

no no don't feel like an idiot!

OpenStudy (anonymous):

We've all had to struggle through this process, even the smartest people on here.

OpenStudy (anonymous):

Although elegant, that approach can be confusing for first-timers.

OpenStudy (anonymous):

sorry, not trying to offend :-)

OpenStudy (boxman61):

no offence, ive never liked algebra... especially when the given examples are not even close thank you all again

OpenStudy (anonymous):

@luis It came out at rude and I apologize for that. Very sorry

OpenStudy (phi):

btw Luis, in your first post you flipped a minus sign. ooops (do it myself sometimes)

OpenStudy (anonymous):

@Boxman61 when you start out, it's easiest to do things in more steps rather than less so that your mind can work through the whole process. I hope your instructor feels the same way as I do.

OpenStudy (boxman61):

her husband wrote the book so i doubt it...

OpenStudy (anonymous):

Oh boy! I wouldn't even wanna got here with her, lol. Have you sought out Google for learning topics in algebra? A lot of times, I can google part of a story problem and something very similar will pop up. The numbers are just different. I don't encourage cheating, but I think it's step to actually learning the process (I mean if the book's not helping then what the heck. You'll learn it eventually anyway).

OpenStudy (boxman61):

i use it when needed... dont even like using OpenStudy but sometimes this book makes no sense

OpenStudy (anonymous):

Have you considered using Paul's online notes as an extra resource?

OpenStudy (anonymous):

This instructor is God imho http://tutorial.math.lamar.edu/Classes/Alg/Alg.aspx